Circular No. 7540

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fd IN UGC 291 T. Puckett, Mountain Town, GA; and S. Gauthier, London, ON, re-

port the discovery of an apparent supernova (mag 17.7) on an unfiltered CCD frame (limiting mag 20.2) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Dec. 8.09 UT. SN 2000fd is located at α = 0^{h}29^{m}12^{s}.09, δ = +33^{o}06′29′′.5 (equinox 2000.0), which is 7′′.3 west and 14′′.8 north of the center of UGC 291. The new object was also present and confirmed on unfiltered CCD frames taken on Dec. 8.97. The candidate does not appear on unfiltered images taken by Puckett during July 20–Oct. 22, and it does not appear on Palomar Sky Survey images taken on 1989 Sept. 2 and 3 (limiting mag ∼ 21.0).

SUPERNOVA 2000fe IN UGC 4870 M. Schwartz, Cottage Grove, OR; and W. D. Li and A. V. Filippenko,

University of California at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova with the Tenagra III 0.5-m automated telescope. The object was discovered and confirmed on unfiltered images taken on Dec. 8.4 (mag ∼ 16.6) and 9.3 UT (mag ∼ 16.4). SN 2000fe is located at α = 9^{h}14^{m}55^{s}.15, δ = +46^{o}54′02′′.0 (equinox 2000.0), which is 5′′.1 west and 9′′.5 south of the nucleus of UGC 4870. An image of the field taken with the same telescope on Dec. 5.4 showed nothing at this position (limiting mag ∼ 18.5). SN 1966A also appeared in UGC 4870.

COMET C/1999 T1 (McNAUGHT-HARTLEY) Total visual magnitude estimates: Nov. 23.29 UT, 8.4 (A. Amorim,

Florianopolis, Brazil, 0.14-m reflector); 26.72, 8.5 (M. Mattiazzo, Wallaroo, S. Australia, 25×100 binoculars); 28.84, 8.1 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector); Dec. 2.83, 8.4 (A. Pearce, Nedlands, W. Australia, 0.20-m reflector); 4.50, 8.6 (A. Hale, Cloudcroft, NM, 11×80 binoculars); 5.83, 7.8 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector).

COMET 41P/TUTTLE-GIACOBINI-KRESÁK Corrigendum. On IAUC 7536, last line, for (M. Tsumura, Wakayama,

Japan, 0.32-m reflector). read (Nagai).

2000 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green